# Talking Wikidata: Communication patterns and their impact on community engagement in collaborative knowledge graphs

### v1

@@ 6.2.1 Summary
6.3 RQ3 - Do discussions impact the editor’s engagement?
The study aimed to investigate the factors that influence whether a Wikidata editor could be active. In this case, the ablation study was focused on editors and included editor’s features (EF), text embedding for all their posts (Tavr), and graph embedding for the editor’s topology (Geditor) (see Figure 4).
Table 5 presents the experiment results. Among the features analysed, account age showed the best performance. This may suggested that the longer an editor has been active on Wikidata, the more likely they were to remain active. Upon examining the status of editors with more than six years of account age, we found that $85\%$ of them were active. We further investigated this hypothesis by testing the statistical difference between the account age and the status of editors using the Mann-Whitney U non-parametric test [29]. The results showed that there was a significant difference between the account age of active and inactive editors, with all the values of account age for the active editors being greater than those of the inactive editors (p-value = 0).
It is worth noting that the rest of editors’ features, and specifically EFaccess level, presented very low influence on their status. This means that editors with higher access levels were not likely to be active compared to basic level users.
We also discovered that editor’s text embedding also showed an influence on their status. This text included all the posts that an editor wrote. This result showed that discussions had an effect on editors’ engagement. Finally, we found that a combination of the editor’s account age and the content of their posts produced the best performance.
Table 5: Results for the model predicting the editors’ status.
@@ 7 Discussion
We analysed the Wikidata discussions by studying the structure of editors’ interactions network and exploring what influences the continuity of a discussion and editors’ engagement. Our findings highlight the value of investigating communication interactions to understand what influences engagement and support collaborative communities to improve practice and retain and increase the number of members and their contributions.
@@
We found that Wikidata discussion interactions form a small world network. The small world characteristics facilitate the propagation of information much faster than random networks [25, 44]. Wikidata could benefit from this for tool adoption, spam detection or vandalism. However, this quick spread of information could potentially be harmful in case of deliberate spread of “fake news”, e.g. fake practices.
@@
Table 6 presents network metrics for four online communities, including Wikidata, forming small world networks. Comparing the network metrics we observe that Wikidata is more similar to Wikipedia than to MSN or Slashdot. A high cluster coefficient indicates a closely connected community where individuals are connected to their friends, and their friends are also connected to each other. It is worth noting that the Wikidata network was found to have an average cluster coefficient of six and $12$ times higher than the MSN and the Slashdot community, respectively, indicating a stronger sense of cohesion and community, similar to Wikipedia. Additionally, the Wikidata network has the lowest average and longest shortest path compared to the other three communities, which means that the nodes in the Wikidata network are closer connected to each other, making it easier and faster to traverse from one node to another. Furthermore, unlike Slashdot, Wikidata forms a disassortative network. This means that ‘popular’ editors tend to communicate with ‘less popular’ editors within the Wikidata community. However, Wikidata seems to demonstrate different assortativeity from the Wikipedia user network, which has neutral assortativity but presents similarities with the Wikipedia Wikiprojects.
@@
The closely connected community for Wikidata can be explained from the variety of topics/channels i.e. property proposal, item talk pages, etc. Many editors are not limited to one kind of discussion channel but converse to different channels about different issues, technical or administrative. This behaviour may enhance the connection of editors with different responsibilities in the community. This system is similar to Wikipedia, which may be why the two projects present more similarities than the other two communities. However, degree assortativity is significantly different for Wikidata, showing the commitment of ‘popular’ editors to support less ‘popular’ members which is likely to be found also in Wikipedia Wikiprojects.
@@
We found that the main influence for a discussion to receive a response is its content. Our prior work suggested that reasons for no responses in Wikidata could be the numerous item talk pages leaving posts unattended and the lack of guidance for using talk pages compared to Wikipedia [22]. In that study, we explored topics of item discussions showing that $46\%$ of topics were related to knowledge engineering. The editors discussed content page management, suggested or reported changes, and asked for help and specifications. The complexity of the subject, particularly for a group of people with diverse KE experience, may made it difficult for editors to express the subjects in a simple and informative way. This may prevented responses to obscure posts.
@@
The complexity of KE topics suggest that guidance in asking questions and discussing changes may increase the possibility of continuing a conversation. This may be possible with the use of templates for specific cases. Wikidata already use templates for specific discussion channels like Request for a comment and (Property proposal) to help the community receive all the information needed to make a decision on a subject. The use of simple templates also for other discussion pages for questions and requests may enhance communication and help, especially new members, to receive answers. Guidance on discussions has been emphasised previously for the GitHub community to support members’ participation [15].
In addition, another suggestion to reduce the posts with no responses is for Wikidata to adopt a system similar to the annotation system for talk pages based on Semantic Web technologies suggested for Wikipedia [39]. The authors suggested a lightweight annotation process, similar to other projects like HyperNews, to help classify the type of edits that happen in the content pages. However, a similar system could be used by Wikidata to classify the posts to help, question, idea, warning etc., to help the community monitor the numerous posts in talk pages, and enhance attention special to new members.
@@
Our findings suggested that editors’ engagement was influenced by their account age and the content of their discussions. We show that the longer the account age, the more likely to be active. This indicated that one important factor for engagement was the first period an editor joined the community. Our previous study in Wikidata discussions [22], emphasised a high number of posts with no response, $50\%$ of posts in item talk pages. Furthermore, in this study, the analysis revealed the influence of no responses to editors’ engagement with $47\%$ of editors who did not receive a response to their posts dropped out, and the rest did not attempt to participate in any discussion in the future. This emphasises the importance of our previous suggestions to create a system to reduce posts with no responses, not only for a conversation to continue but also for member engagement, especially for new members of the community.
@@
In addition, another factor influencing the editors’ engagement was the content of their posts. In the Wikidata scenario, we previously showed that Wikidata item talk pages had low frequency in conflicts [22]. This could indicate that editors did not receive aggressive or confrontational feedback leading to drop out, but the posts’ character, e.g. most posts are questions, may influenced the prediction. This is different from open source software where Filippova et al. showed that some types of conflicts impacted the intention of members to email in the project [10]. As mentioned earlier, in our previous study, we argued that the topics discussed mainly relate to KE activities. The complexity of KE, especially for members with no prior experience in building KGs, with questions receiving insufficient or confusing feedback could lead to disappointment and resignation. These results highlight again the importance of first experiences for engagement when new members join the community. The community may need to monitor whether new members ask for help, developing practices for guidance and support to encourage them so as to overcome a period of time participating to stay engaged.
@@ 7.1 Limitations
For this study, we used a sample of discussion and only quantitative analysis to investigate the impact of text on the course of discussion and editors’ engagement. However, qualitative analysis similar to previous studies about the content discussed [22, 19] could reveal insights into the influential topics. Furthermore, we investigated the influential factors only in item talk pages. In contrast, including different types of discussion channels may provide extra details to the prediction models. Finally, we tested one network structure for the predictions. Experimenting with different designs may improve results for the post and editors’ topology, giving another understanding of these predictions.
@@ 8 Conclusion
We investigated the Wikidata editors’ interactions through their lens of discussions. We used a mixed method approach using descriptive statistics, network analysis, and graph and text embedding vectors. We found that the Wikidata community forms a small world network with news propagating fast across the network. At the same, Wikidata ‘popular’ editors seem to communicate a lot with less ‘popular’ editors, a behaviour similar to Wikipedia Wikiprojetcs. Additionally, we suggested that the longer an editor stays in the community and feels inclusive, the more likely to stay engaged in the project. A suggestion could be for Wikidata to control the numerous comments with no responses in order to support particularly new editors and make them feel included. We hope our findings can support Wikidata in developing new practices to help new members find their steps in the project to stay engaged, as well as help project designers transmit new tactics to the community. Future work could investigate the content of the text and justify what makes a post a good candidate to receive an answer and what discourages members from responding to others. In the same vein, it would be interesting to investigate the text that makes editors stop contributing. This will give an in-depth understanding and enable Wikidata to improve practices and tools.

### v2

@@ 6.2.1 Summary
6.3 RQ3 - Do discussions affect the editor’s engagement?
The study aimed to investigate the factors that influence whether a Wikidata editor could be active. In this case, the ablation study was focused on editors and the features of editors (EF), text embedding for all their posts (Tavr), and graph embedding for the topology of editors (Geditor) (Table 4).
Table 6 presents the experiment results. Similar to the results for predicting a response, the results for predicting the status of editors showed a balance between precision, recall, and F1 scores for the combination of features (e.g., EF + Geditor), suggesting a balanced and high-performing model. In contrast, in the cases of individual features for the characteristics of editors (e.g., EFaccess level), the model tended to make more positive predictions than false positives to all features except for EFage. This suggests that we cannot predict whether an editor is active for sure using only one characteristic. However, we can see that the features for editors, and specifically EFaccess level, presented very low scores. These results indicate that editors with higher access levels were not likely to be active compared to basic level users.
Table 6 shows that among the features analysed, account age showed the best performance, at $0.72\%$. This may suggest that the longer an editor had been active on Wikidata, the more likely they were to remain active. Upon examining the status of editors with more than six years of account age, we found that $85\%$ of them were active. We further investigated this hypothesis by testing the statistical difference between the account age and the status of editors using the Mann-Whitney U non-parametric test [29]. The results showed that there was a significant difference between the account age of active and inactive editors, with all the values of account age for the active editors being greater than those of the inactive editors (p-value = 0).
We also discovered that text embedding of editors had an influence on their status, with $0.69\%$ accuracy. This text included all the posts that an editor wrote. This result showed that discussions affect the engagement of editors. Additionally, the combination of account age and the content of posts produced the best performance for the status of editors, with $0.75\%$ accuracy.
Table 6: Performance results of the ablation study for the features individually and in combination. The ML model predicted whether an editor was active.
@@ 7 Discussion
We analysed the Wikidata discussions by studying the structure of editors’ communication networks and exploring what affects the continuity of a discussion and the engagement of editors. Our findings highlight the value of investigating communication interactions to understand what influences engagement and support collaborative communities to improve practice and retain and increase the number of members and their contributions. In this vein, based on our findings, we have developed the following list of recommendations for Wikidata and semantic web communities, which we further elaborate on in this section:
  • Incorporate instructions and tutorials on how to use talk pages and specialised discussion channels
  • Develop a post-monitor system similar to HyperNews and GitHub
  • Develop a mentoring system between long-lasting and new members
  • Incorporate templates to assist editors in discussing complex KE topics similar to GitHub
RQ1 We found that a high number of editors had no edges, meaning that they did not receive responses to their posts. We showed that the response to a post was related to the status of editors, with $47\%$ of editors who did not receive responses to their posts being inactive. Additionally, our previous study on the content of Wikidata discussions [22] showed that $50\%$ of threads in item talk pages were posts without receiving responses, which could potentially impact the engagement of members, particularly the new ones. To support new members and improve practices, we previously suggested the creation of clear instructions on how to use discussions in Wikidata [22]. Another suggestion to reduce the number of posts with no responses is the development of a system to monitor posts. An annotation system was previously suggested for Wikipedia, similar to other projects like HyperNews [39]. This could help classify the type of edits that happen in the content pages. A similar system could be used by Wikidata to classify the posts to help, question, idea, warning and so on, and thus bring them to people’s attention. This system would allow the community to keep track of posts on talk pages and provide special attention to new members.
@@
We found that Wikidata discussion interactions form a small world network. The small world characteristics: facilitate the propagation of information much faster than random networks; suggests a robust network resilient to node removal, but yet vulnerable to targeted attacks; and benefit scalability, that is, maintaining the small world characteristics even as networks grow in size [25, 44, 52]. Wikidata could benefit from this to develop strategic planning for tool adoption, or the detection of spam or vandalism. However, the community should be aware of the quick spread of information because this could potentially be harmful in case of deliberate spread of “fake news”, such as fake practices.
@@
Table 7 shows network metrics for four online communities, including Wikidata, forming small world networks. Comparing the network metrics we note that Wikidata is more similar to Wikipedia than to MSN or Slashdot. A high clustering coefficient indicates a closely connected community where individuals are connected to their friends, and their friends are also connected to each other. It is worth mentioning that the Wikidata network was found to have an average cluster coefficient of six and $12$ times higher than the MSN and the Slashdot community, respectively, indicating a stronger sense of cohesion and community, similar to Wikipedia. Additionally, the Wikidata network has the lowest average and longest shortest path compared to the other three communities, which means that the nodes in the Wikidata network are closer connected to each other, making it easier and faster to traverse from one node to another. Furthermore, unlike Slashdot and Wikipedia, Wikidata formed a disassortative network. This means that well-connected editors tended to communicate with low-connected editors within the Wikidata community. This indicates that Wikidata was an inclusive community, more integratively connected than the other two, while the other two had random connections between nodes. This is similar to Wikipedia Wikiprojects, where we find small groups of editors closely collaborating.
@@
The closely connected community for Wikidata could be explained from the variety of discussion channels, such as property proposal, administrators’ noticeboard, and item talk pages. Some editors might not be limited to one kind of discussion channel, but might rather converse and share opinions to to different channels about different issues, technical or administrative. Alternatively, the majority of editors might tend to participate in general issues discussion channels like request for comments and subsequently to be connected. This behaviour may enhance the connection of editors with different responsibilities in the community. Although Wikipedia shares a similar system with various discussion channels, its seems that the two projects have different community dynamics. Wikipedia community presents random connections, while Wikidata is an inclusive community with strongly connected groups. This may suggest that Wikipedia members may tend to use the article talk pages to communicate more than other discussion channels, while in Wikidata we found a low use of item talk pages [22]
We found that $20\%$ of editors made more than 10 posts, suggesting that they made the most contributions in discussions and they were potentially responsible for key decisions. This is more balanced than the schema.org community [20] and the $90-9-1$ rule for participation inequality in social media and online communities [90_9_1], which suggests that $90\%$ of members are lurkers who never contribute, $9\%$ contribute a little, and $1\%$ account for almost all the activity. This suggests that in Wikidata, there is a higher proportion of editors participating at a low but significant level. In addition, results showed that $10\%$ of editors had been members of the community for more than $10$ years, which is consistent since the earlier stages of the project [38]. This indicates that the Wikidata community has a group of committed and experienced members. Leveraging their expertise, these members could be advised to guide the development of new systems to improve practices, such as recommendation systems. Furthermore, in the case of creating a system to classify posts to enhance monitoring and limit posts without responses, these experienced members could help by responding to posts and serving as mentors to new members. The participation of core contributors in discussions and mentoring of new members has been also emphasised by Hata et al. [15] for the GitHub projects.
@@
RQ2 We found that the primary factor influencing whether a discussion received a response was its content. This encompassed the topic and the clarity in defining it. Our prior work [22] suggested that reasons for no responses in Wikidata could be the numerous items which may leave their talk pages unattended and the lack of guidance for using talk pages compared to Wikipedia. In this study [22], we explored topics of item discussions, showing that $46\%$ of topics were related to KE. The editors discussed content page management, suggested or reported changes, and asked for help and specifications related to Wikidata tasks. The complexity of the subject, particularly for a group of people with diverse KE experience, may make it difficult to express the subjects in the simplest and most informative way. This may mean that not everyone in the community feels that they have the expertise to respond. This argument is enhanced by the findings that the access level of an editor did not have an influence on a post to receive a response. This indicates that even posts from editors with higher editing rights and some experience in KE may remain unanswered.
@@
The complexity of KE topics suggests that guidance in asking questions and discussing changes may increase the possibility of continuing a conversation. This may be possible with the use of templates for specific cases. Wikidata already use templates for specific discussion channels like request for a comment and (property proposal) to help the community receive all the information needed to make a decision on a subject. The use of templates for other discussion pages for questions and requests may enhance communication and help new members, in particular, to receive answers. Guidance on discussions has been emphasised previously for the GitHub community to support the participation of members [15].
@@
RQ3 Our findings suggest that the engagement of editors is influenced by their account age and the content of their discussions. We showed that the higher the account age, the more likely an editor was to be active. This indicates that one important factor for engagement is the time when an editor first joins the community. Our previous findings from $RQ2$ showed that the status of editors was related to whether their posts received responses, with $47\%$ of editors who did not receive a response to their posts dropping out of Wikidata, and the rest not attempting to participate in any discussion in the future. These results highlight the influence of no responses on the engagement of editors and emphasise the importance of our previous suggestions to create a system to reduce posts with no responses, especially for item talk pages, which account for $50\%$ of posts did not receive a response [22]. This could improve both the continuation of conversations and the engagement of members in activities, primarily for new members of the community.
@@
In addition, another factor influencing the engagement of editors was the content of their posts. We previously showed that Wikidata item talk pages had low frequency in conflicts ($7-9\%$ disagreements and up to $1\%$ intensified conflict) [22]. This could indicate that editors do not receive aggressive or confrontational feedback leading to drop out, but the posts’ character, such as the fact that many posts are questions, may have influenced the prediction of editors’ status. This is different from open source software where Filippova et al. [10] showed that some types of conflicts impacted the intention of members to email in the project. As mentioned earlier, in our previous study, we argued that the topics discussed mainly relate to KE activities [22]. The complexity of KE, especially for members with no prior experience in building KGs, with questions receiving insufficient or confusing feedback, could lead to disappointment and drop outs. These results highlight again the importance of first experiences related to engagement when new members join the community. The community may need to monitor whether new members ask for help, developing practices for guidance and support to encourage them to stay engaged. Based on previous findings, we recommended creating an annotation system to monitor posts in order to decrease the number of posts without responses. This system could use advanced technology like generative AI to provide better support for new members, helping them to receive answers and encouraging their continued contributions.
It is worth noting that editor characteristics, such as the access level, did not influence either the communication or the participation of editors. For posts receiving a response, this was expected considering previous findings in $RQ1$, with basic level users being responsible for a greater share of contributions in discussions.
@@ 7.1 Limitations
For the network analysis, due to the considerable size of the full network, in some cases, we used a smaller random sample to be able to calculate specific metrics. However, the sampling could corrupt the structural properties of the network and result in inconsistent findings. Being able to manage the full size of the network could lead to more robust patterns for metrics like shortest paths and provide extra information with metrics like homophily. Additionally, in our analysis, we assumed that if two editors took part in the same discussion thread, they were connected. However, in threads with a high number of posts, some editors might participate in the same topic of discussion without directly addressing each other. Alternatively, a future study could track the replies to posts in order to connect editors based on who is responding to whom.
For the prediction models, we used a sample of discussion and only quantitative analysis to investigate the impact of text on the course of discussion and the engagement of editors. However, qualitative analysis similar to previous studies about the content discussed [22, 19] could reveal insights into the influential topics. Furthermore, we investigated the influential factors only in item talk pages. In contrast, including different types of discussion channels may provide extra details to the prediction models. Finally, we tested one network structure for the predictions. Experimenting with different designs, such as connecting posts with reply direction, may improve results for the post and editors’ topology, giving another understanding of these predictions.
@@ 8 Conclusion
We investigated the Wikidata editors’ interactions through their lens of discussions. We used a mixed method approach using descriptive statistics, network analysis, and graph and text embedding vectors. We found that the Wikidata community forms a small world network similar to MSN, Slashdot, and Wikipedia. However, Wikidata presents stronger connections than the other communities. Our findings suggest that Wikidata is a valuable example for other communities to mimic its practices in order to build collaborative communities that are inclusive and resilient to dropouts. Wikidata ‘popular’ editors seem to communicate a lot with less ‘popular’ editors which shows a significant difference between Wikipedia and Wikidata communities. Despite their numerous similarities in terms of the MediaWiki platform, administrative support, and moderation tools, the two communities present differences in communication and connectivity, with Wikidata editors having more connections to each other. Additionally, we suggested a list of practices to improve communication and engagement with clear instructions about discussions, a post-monitor system, a mentoring system, and communication templates. The recommendation list can help Wikidata and collaborative communities in developing new practices to help new members find their steps in the project to stay engaged, as well as help project designers transmit new tactics to the community.
Beyond Wikidata, our study can provide insights and guidance to semantic web communities on how to build and optimise collaborative ecosystems for sustainability, growth, and quality improvement. Our findings on how Wikidata is connected and what affects communication and participation can be used as a guide not only for general domain KGs and ontologies, but also in specialised domains like healthcare, cultural heritage, and open data initiatives, such as government or public-sector linked data projects. Our suggested framework can be used as a base for future studies. Other communities can use this framework and extend it by adding extra specialised features based on the community peculiarities, as well as features on how cultural and linguistic diversity affects communication and participation in global communities. In addition, a future analysis can utilise the framework and the dataset to run the analysis again in order to validate community evolution over time and perhaps after adapting new practices and improvements based on our recommendations. The framework and dataset can be used to deepen the analyses on the topics affecting communication and participation and identify discussions for qualitative analysis or sentiment analysis. A further analysis on the content of the text can justify what makes a post a good candidate to receive an answer and what discourages members from responding to others. In the same vein, it would be interesting to investigate the text that makes editors stop contributing. Furthermore, by linking the dataset discussions to the parallel entity activity, a future study could explore how discussions influence productivity. Finally, interviews with Wikidata editors could add valuable insights about their experience in the project, their collaborations, and ideas for improvements.